The following data are accumulated by Eco Labs, Inc. in evaluating two competing capital investment proposals:

Testing Equipment Vehicle
Amount of investment $68,000 $92,000
Useful life 4 years 7 years
Estimated residual value 0 0
Estimated total income over the useful life $6,800 $30,590

Determine the expected average rate of return for each proposal. If required, round your answers to one decimal place.

Testing Equipment %
Vehicle %
Answer #1

Expected average rate of return = Average income/Average investment

testing equipment = (6800/4)/34000 = 5%

Vehicle = (30590/7)/46000 = 9.5%